What two number when added equal 13 and when subtracted equal 43
Answer by LinnW(1048)

x + y = 13
x - y = 43
Add the two equations
2x = 56
divide each side by 2
x = 28
substitute 28 for x in x + y = 13
28 + y = 13
subtract 28 from each side
y = -15
So the numbers are -15 and 28